# Two-grid convergence theory for symmetric positive semidefinite linear systems

### v1

@@ 4.1. The linear case
(4.12a) $\lambda_{\min}\big{(}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{)}$ $=0,$
(4.12b) $\lambda_{\max}\big{(}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{)}$ $=1-\delta_{\rm TG},$
@@
The identity (4.12a) is obvious, because $\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\succeq 0$ is singular. Next, we focus on the proof of (4.12b).
Since $A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\succeq 0$ and $\lambda\big{(}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big{)}\subset[0,1]$, it follows from (3.10) that $X_{1}\succeq 0$ and $\lambda(X_{1})\subset[0,1]$. By (3.9) and (3.10), we have
@@
We then prove that $\operatorname*{rank}(X_{1})=s$ (i.e., $X_{1}\succ 0$) if and only if
@@
  • Recall that
    $\mathcal{R}(\Pi)=\mathcal{R}\big{(}A^{\frac{1}{2}}P\big{)}\quad\text{and}\quad\mathcal{N}(\Pi)=\mathcal{N}\big{(}P^{T}A^{\frac{1}{2}}\big{)}.$
    If (4.16) holds, then
@@
    which, together with (4.15), leads to $\operatorname*{rank}(X_{1})=s$.
@@
    which follows from the relations $\mathcal{N}(\Pi)\subseteq\mathcal{N}\big{(}\Pi A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\Pi\big{)}$ and
@@
Combining (4.14), (4.17), and (4.18), we can arrive at the identity (4.12b). ∎
@@
The relation (4.16) will hold if $\widetilde{M}\succeq 0$ and $\mathcal{N}(\widetilde{M})\cap\mathcal{R}(A)=\{0\}$, since
$\mathcal{N}\big{(}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big{)}=\mathcal{N}\big{(}\widetilde{M}^{\frac{1}{2}}A^{\frac{1}{2}}\big{)}=\mathcal{N}\big{(}\widetilde{M}A^{\frac{1}{2}}\big{)}=\mathcal{N}(A)\quad\text{and}\quad\mathcal{R}\big{(}A^{\frac{1}{2}}P\big{)}\subseteq\mathcal{R}(A).$
Therefore, a sufficient condition for (4.16) to hold is $\widetilde{M}\succ 0$, which can be satisfied by the weighted Jacobi and Gauss–Seidel smoothers, as discussed in Remark 2.6.
We are now ready to present a two-sided estimate for $\|E_{\rm ITG}\|_{A}$.
@@
Under the assumptions of Algorithm 2 and Lemma 4.2, the inexact two-grid convergence factor $\|E_{\rm ITG}\|_{A}$ satisfies
(4.19) $\mathscr{L}(\beta_{2})\leq\|E_{\rm ITG}\|_{A}\leq\mathscr{U}(\beta_{1})$
with
$\mathscr{L}(\beta_{2})=\sqrt{1-\min\big{\{}\lambda_{n-r+1}(\widetilde{M}A)+\beta_{2}(1-\delta_{\rm TG}),\,\sigma_{\rm TG}\big{\}}}$
@@
(4.20) $\mathscr{U}(\beta_{1})=\sqrt{1-\max\big{\{}\lambda_{n-r+1}(\widetilde{M}A),\,\beta_{1}\sigma_{\rm TG},\,\sigma_{\rm TG}-(1-\beta_{1})(1-\delta_{\rm TG})\big{\}}},$
where $\widetilde{M}$, $\sigma_{\rm TG}$, $\beta_{1}$, $\beta_{2}$, and $\delta_{\rm TG}$ are given by (3.5), (3.8), (4.6a), (4.6b), and (4.13), respectively.
@@
In view of (4.10), one can get a two-sided estimate for $\|E_{\rm ITG}\|_{A}$ by bounding the eigenvalues $\lambda_{n-r+1}(F_{1})$ and $\lambda_{n-r+1}(F_{2})$.
A lower bound for $\lambda_{n-r+1}(F_{1})$: By (4.11a), we have
$\lambda_{n-r+1}(F_{1})$ $\geq\lambda_{n-r+1}\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}+\beta_{1}\lambda_{1}\big{(}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{)}$
$=\lambda_{n-r+1}(\widetilde{M}A),$
where we have used the facts (4.12a) and $\lambda\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}=\lambda\big{(}A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big{)}$. In addition, we have
$\lambda_{n-r+1}(F_{1})$ $=\lambda_{n-r+1}\big{(}(1-\beta_{1})A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}+\beta_{1}F_{\rm TG}\big{)}$
$\geq(1-\beta_{1})\lambda_{1}\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}+\beta_{1}\lambda_{n-r+1}(F_{\rm TG})$
$=\beta_{1}\sigma_{\rm TG}$
and
$\lambda_{n-r+1}(F_{1})$ $=\lambda_{n-r+1}\big{(}F_{\rm TG}-(1-\beta_{1})\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{)}$
$\geq\lambda_{n-r+1}(F_{\rm TG})-(1-\beta_{1})\lambda_{n}\big{(}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{)}$
$=\sigma_{\rm TG}-(1-\beta_{1})(1-\delta_{\rm TG}),$
where, in the last equality, we have used the fact (4.12b). Hence,
(4.21) $\lambda_{n-r+1}(F_{1})\geq\max\big{\{}\lambda_{n-r+1}(\widetilde{M}A),\,\beta_{1}\sigma_{\rm TG},\,\sigma_{\rm TG}-(1-\beta_{1})(1-\delta_{\rm TG})\big{\}}.$
An upper bound for $\lambda_{n-r+1}(F_{2})$: Using (4.11b) and (4.12b), we obtain
$\lambda_{n-r+1}(F_{2})$ $=\lambda_{n-r+1}\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}+\beta_{2}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{)}$
$\leq\lambda_{n-r+1}\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}+\beta_{2}\lambda_{n}\big{(}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{)}$
$=\lambda_{n-r+1}(\widetilde{M}A)+\beta_{2}(1-\delta_{\rm TG}).$
Moreover,
$\lambda_{n-r+1}(F_{2})$ $=\lambda_{n-r+1}\big{(}F_{\rm TG}-(1-\beta_{2})\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{)}$
$\leq\lambda_{n-r+1}(F_{\rm TG})-(1-\beta_{2})\lambda_{1}\big{(}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{)}$
$=\sigma_{\rm TG},$
where, in the last equality, we have used the fact (4.12a). Thus,
(4.22) $\lambda_{n-r+1}(F_{2})\leq\min\big{\{}\lambda_{n-r+1}(\widetilde{M}A)+\beta_{2}(1-\delta_{\rm TG}),\,\sigma_{\rm TG}\big{\}}.$
The estimate (4.19) then follows by combining (4.10), (4.21), and (4.22). ∎
@@
By (4.11a) and (4.12b), we have
$1-\delta_{\rm TG}$ $=\lambda_{n}\big{(}\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}\big{)}$
$=\lambda_{n}\big{(}F_{\rm TG}-A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}$
$\geq\lambda_{n-r+1}(F_{\rm TG})-\lambda_{n-r+1}\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}$
$=\sigma_{\rm TG}-\lambda_{n-r+1}(\widetilde{M}A),$
which gives
$\sigma_{\rm TG}\leq 1-\delta_{\rm TG}+\lambda_{n-r+1}(\widetilde{M}A).$
According to (4.6a) and (4.6b), we deduce that $0<\beta_{1}\leq\beta_{2}\leq 1$.
  • (i)
    If $\beta_{2}\rightarrow 0^{+}$, then $F_{i}\rightarrow A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}$ for $i=1,2$, which, together with (4.10), yields
    $\|E_{\rm ITG}\|_{A}\rightarrow\sqrt{1-\lambda_{n-r+1}\big{(}A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big{)}}=\sqrt{1-\lambda_{n-r+1}(\widetilde{M}A)}.$
    This asymptotic behavior can also be observed from (4.19) as $\beta_{2}\rightarrow 0^{+}$.
  • (ii)
    In the special case $B_{\rm c}=A_{\rm c}$ (i.e., $\beta_{1}=\beta_{2}=1$), we have
    $\mathscr{L}(1)=\mathscr{U}(1)=\sqrt{1-\sigma_{\rm TG}},$
    which leads to the convergence identity (3.7).
@@
4.2. The general case
In this subsection, we present a convergence estimate for Algorithm 2 with general coarse solver $\mathscr{B}_{\rm c}\llbracket\cdot\rrbracket$, as described in the following theorem.
Theorem 4.10.
Under the assumptions of Algorithm 2, if
(4.23) $\big{\|}\mathbf{e}_{\rm c}-\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket\big{\|}_{A_{\rm c}}\leq\varepsilon\|\mathbf{e}_{\rm c}\|_{A_{\rm c}}$
for some $\varepsilon\in[0,1)$, then
(4.24) $\|\mathbf{u}-\mathbf{u}_{\rm ITG}\|_{A}\leq\mathscr{U}(1-\varepsilon^{2})\big{\|}\mathbf{u}-\mathbf{u}^{(0)}\big{\|}_{A},$
where $\mathscr{U}(\cdot)$ is given by (4.20).
Proof.
The condition (4.23) reads
$2\big{(}A_{\rm c}A_{\rm c}^{\dagger}\mathbf{r}_{\rm c}\big{)}^{T}\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket\geq(1-\varepsilon^{2})\|\mathbf{e}_{\rm c}\|_{A_{\rm c}}^{2}+\big{\|}\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket\big{\|}_{A_{\rm c}}^{2}.$
Note that $A_{\rm c}A_{\rm c}^{\dagger}$ is a projector onto $\mathcal{R}(A_{\rm c})$ and
$\mathbf{r}_{\rm c}=P^{T}A\big{(}\mathbf{u}-\mathbf{u}^{(1)}\big{)}\in\mathcal{R}\big{(}P^{T}A^{\frac{1}{2}}\big{)}=\mathcal{R}(A_{\rm c}).$
We then have
(4.25) $2\mathbf{r}_{\rm c}^{T}\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket\geq(1-\varepsilon^{2})\|\mathbf{e}_{\rm c}\|_{A_{\rm c}}^{2}+\big{\|}\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket\big{\|}_{A_{\rm c}}^{2}.$
From Algorithm 2, we have
$\mathbf{u}-\mathbf{u}_{\rm ITG}=\mathbf{u}-\mathbf{u}^{(1)}-P\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket.$
Then
$\|\mathbf{u}-\mathbf{u}_{\rm ITG}\|_{A}^{2}$ $=\big{(}\mathbf{u}-\mathbf{u}^{(1)}-P\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket\big{)}^{T}A\big{(}\mathbf{u}-\mathbf{u}^{(1)}-P\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket\big{)}$
$=\big{\|}\mathbf{u}-\mathbf{u}^{(1)}\big{\|}_{A}^{2}-2\big{(}\mathbf{u}-\mathbf{u}^{(1)}\big{)}^{T}AP\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket+\big{\|}\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket\big{\|}_{A_{\rm c}}^{2}$
$=\big{\|}\mathbf{u}-\mathbf{u}^{(1)}\big{\|}_{A}^{2}-2\mathbf{r}_{\rm c}^{T}\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket+\big{\|}\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket\big{\|}_{A_{\rm c}}^{2}$
$\leq\big{\|}\mathbf{u}-\mathbf{u}^{(1)}\big{\|}_{A}^{2}-(1-\varepsilon^{2})\|\mathbf{e}_{\rm c}\|_{A_{\rm c}}^{2}$
$=\big{(}\mathbf{u}-\mathbf{u}^{(1)}\big{)}^{T}A\big{(}\mathbf{u}-\mathbf{u}^{(1)}\big{)}-(1-\varepsilon^{2})\mathbf{r}_{\rm c}^{T}A_{\rm c}^{\dagger}\mathbf{r}_{\rm c}$
$=\big{(}\mathbf{u}-\mathbf{u}^{(1)}\big{)}^{T}A\big{(}\mathbf{u}-\mathbf{u}^{(1)}\big{)}-(1-\varepsilon^{2})\big{(}\mathbf{u}-\mathbf{u}^{(1)}\big{)}^{T}A^{\frac{1}{2}}\Pi A^{\frac{1}{2}}\big{(}\mathbf{u}-\mathbf{u}^{(1)}\big{)}$
$=\big{(}\mathbf{u}-\mathbf{u}^{(1)}\big{)}^{T}A^{\frac{1}{2}}\big{(}I-(1-\varepsilon^{2})\Pi\big{)}A^{\frac{1}{2}}\big{(}\mathbf{u}-\mathbf{u}^{(1)}\big{)},$
where we have used the inequality (4.25). Due to
$\mathbf{u}-\mathbf{u}^{(1)}=(I-MA)\big{(}\mathbf{u}-\mathbf{u}^{(0)}\big{)}$
and
$\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}=I-A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}},$
it follows that
$\|\mathbf{u}-\mathbf{u}_{\rm ITG}\|_{A}^{2}\leq\big{(}\mathbf{u}-\mathbf{u}^{(0)}\big{)}^{T}A^{\frac{1}{2}}(I-F)A^{\frac{1}{2}}\big{(}\mathbf{u}-\mathbf{u}^{(0)}\big{)}$
with
$F=A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}+(1-\varepsilon^{2})\big{(}I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big{)}\Pi\big{(}I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big{)}.$
Then
$\|\mathbf{u}-\mathbf{u}_{\rm ITG}\|_{A}^{2}\leq\bigg{(}1-\min_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\mathbf{v}^{T}F\mathbf{v}}{\mathbf{v}^{T}\mathbf{v}}\bigg{)}\big{\|}\mathbf{u}-\mathbf{u}^{(0)}\big{\|}_{A}^{2}.$
Similarly to the proof of Lemma 3.1, one can prove that
$\min_{\mathbf{v}\in\mathcal{R}(A)\backslash\{0\}}\frac{\mathbf{v}^{T}F\mathbf{v}}{\mathbf{v}^{T}\mathbf{v}}=\lambda_{n-r+1}(F).$
Thus,
$\|\mathbf{u}-\mathbf{u}_{\rm ITG}\|_{A}\leq\sqrt{1-\lambda_{n-r+1}(F)}\,\big{\|}\mathbf{u}-\mathbf{u}^{(0)}\big{\|}_{A}.$
The estimate (4.24) can be obtained by bounding $\lambda_{n-r+1}(F)$ from below. Indeed, it holds that
$\lambda_{n-r+1}(F)\geq\max\big{\{}\lambda_{n-r+1}(\widetilde{M}A),\,(1-\varepsilon^{2})\sigma_{\rm TG},\,\sigma_{\rm TG}-\varepsilon^{2}(1-\delta_{\rm TG})\big{\}};$
see the proof of (4.21). This completes the proof. ∎
@@ 4.2. The general case
$\|\mathbf{u}_{\rm TG}-\mathbf{u}_{\rm ITG}\|_{A}=\big{\|}P\big{(}\mathbf{e}_{\rm c}-\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket\big{)}\big{\|}_{A}=\big{\|}\mathbf{e}_{\rm c}-\mathscr{B}_{\rm c}\llbracket\mathbf{r}_{\rm c}\rrbracket\big{\|}_{A_{\rm c}},$

### v2

@@ 4.1. Convergence analysis
(4.1) $\lambda_{\max}\big(\big(I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big)\Pi\big(I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big)\big)=1-\delta_{\rm TG},$
@@
Since $A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\succeq 0$ and $\lambda\big(A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big)\subset[0,1]$, we get from (3.8) that $X_{1}\succeq 0$ and $\lambda(X_{1})\subset[0,1]$. By (3.7) and (3.8), we have
@@
Next, we prove that $\operatorname*{rank}(X_{1})=s$ (i.e., $X_{1}\succ 0$) if and only if
@@
  • Recall that $\mathcal{R}(\Pi)=\mathcal{R}\big(A^{\frac{1}{2}}P\big)$ and $\mathcal{N}(\Pi)=\mathcal{N}\big(P^{T}A^{\frac{1}{2}}\big)$. If (4.5) holds, then
@@
    which, together with (4.4), yields $\operatorname*{rank}(X_{1})=s$.
@@
    which follows from the relations
    $\mathcal{N}(\Pi)\subseteq\mathcal{N}\big(\Pi A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\Pi\big)$
    and
@@
Combining (4.3), (4.6), and (4.7), we can arrive at the identity (4.1). ∎
@@
If $\widetilde{M}\succeq 0$ and $\mathcal{N}(\widetilde{M})\cap\mathcal{R}(A)=\{0\}$, then (4.5) holds, since
$\mathcal{N}\big(A^{\frac{1}{2}}\widetilde{M}A^{\frac{1}{2}}\big)=\mathcal{N}\big(\widetilde{M}^{\frac{1}{2}}A^{\frac{1}{2}}\big)=\mathcal{N}\big(\widetilde{M}A^{\frac{1}{2}}\big)=\mathcal{N}(A)$
and
$\mathcal{R}\big(A^{\frac{1}{2}}P\big)\subseteq\mathcal{R}(A)=\mathcal{N}(A)^{\perp}.$
Obviously, a sufficient condition for (4.5) to be valid is $\widetilde{M}\succ 0$, which can be satisfied by the weighted Jacobi and Gauss–Seidel smoothers, as discussed in Remark 2.7.
We are now ready to present a convergence estimate for Algorithm 2.
@@
Let $\rho,\varepsilon\in[0,1)$ be two parameters. If
(4.8) $\|\mathbf{u}-\mathbf{u}_{\rm TG}\|_{A}\leq\rho\|\mathbf{u}-\mathbf{u}^{(0)}\|_{A}$
@@
(4.9) $\|\mathbf{e}_{\rm c}-\hat{\mathbf{e}}_{\rm c}\|_{A_{\rm c}}\leq\varepsilon\|\mathbf{e}_{\rm c}\|_{A_{\rm c}},$
then
(4.10) $\|\mathbf{u}-\mathbf{u}_{\rm ITG}\|_{A}\leq\sqrt{\rho^{2}+\varepsilon^{2}\big(1-\max\big\{\rho^{2}+\lambda_{n-r+1}(\widetilde{M}A),\,\delta_{\rm TG}\big\}\big)}\|\mathbf{u}-\mathbf{u}^{(0)}\|_{A},$
where $\widetilde{M}$ and $\delta_{\rm TG}$ are given by (3.4) and (4.2), respectively.
@@
The condition (4.9) reads
(4.11) $2\mathbf{r}_{\rm c}^{T}\hat{\mathbf{e}}_{\rm c}\geq(1-\varepsilon^{2})\|\mathbf{e}_{\rm c}\|_{A_{\rm c}}^{2}+\|\hat{\mathbf{e}}_{\rm c}\|_{A_{\rm c}}^{2}.$
From Algorithm 2, we have
$\mathbf{u}-\mathbf{u}_{\rm ITG}=\mathbf{u}-\mathbf{u}^{(1)}-P\hat{\mathbf{e}}_{\rm c}.$
Then
$\|\mathbf{u}-\mathbf{u}_{\rm ITG}\|_{A}^{2}$ $=(\mathbf{u}-\mathbf{u}^{(1)}-P\hat{\mathbf{e}}_{\rm c})^{T}A(\mathbf{u}-\mathbf{u}^{(1)}-P\hat{\mathbf{e}}_{\rm c})$
$=\|\mathbf{u}-\mathbf{u}^{(1)}\|_{A}^{2}-2(\mathbf{u}-\mathbf{u}^{(1)})^{T}AP\hat{\mathbf{e}}_{\rm c}+\|\hat{\mathbf{e}}_{\rm c}\|_{A_{\rm c}}^{2}$
$=\|\mathbf{u}-\mathbf{u}^{(1)}\|_{A}^{2}-2\mathbf{r}_{\rm c}^{T}\hat{\mathbf{e}}_{\rm c}+\|\hat{\mathbf{e}}_{\rm c}\|_{A_{\rm c}}^{2}$
[…]
$=(\mathbf{u}-\mathbf{u}^{(1)})^{T}A(\mathbf{u}-\mathbf{u}^{(1)})-(1-\varepsilon^{2})\mathbf{r}_{\rm c}^{T}A_{\rm c}^{\dagger}\mathbf{r}_{\rm c}$
$=(\mathbf{u}-\mathbf{u}^{(1)})^{T}A(\mathbf{u}-\mathbf{u}^{(1)})-(1-\varepsilon^{2})(\mathbf{u}-\mathbf{u}^{(1)})^{T}A^{\frac{1}{2}}\Pi A^{\frac{1}{2}}(\mathbf{u}-\mathbf{u}^{(1)})$
$=(\mathbf{u}-\mathbf{u}^{(1)})^{T}A^{\frac{1}{2}}\big(I-(1-\varepsilon^{2})\Pi\big)A^{\frac{1}{2}}(\mathbf{u}-\mathbf{u}^{(1)}).$
Due to
$\mathbf{u}-\mathbf{u}^{(1)}=(I-MA)(\mathbf{u}-\mathbf{u}^{(0)}),$
it follows that
(4.12) $\|\mathbf{u}-\mathbf{u}_{\rm ITG}\|_{A}^{2}\leq(\mathbf{u}-\mathbf{u}^{(0)})^{T}\big(E_{1}-(1-\varepsilon^{2})E_{2}\big)(\mathbf{u}-\mathbf{u}^{(0)}),$
where
$E_{1}=(I-MA)^{T}A(I-MA)\quad\text{and}\quad E_{2}=(I-MA)^{T}A^{\frac{1}{2}}\Pi A^{\frac{1}{2}}(I-MA).$
Observe that
$E_{1}\succeq 0,\quad E_{2}\succeq 0,\quad\text{and}\quad E_{1}-E_{2}\succeq 0.$
We next determine upper bounds for $\|\mathbf{u}-\mathbf{u}^{(0)}\|_{E_{1}}^{2}$, $\|\mathbf{u}-\mathbf{u}^{(0)}\|_{E_{2}}^{2}$, and their difference.
  • Since
    $E_{1}=A^{\frac{1}{2}}\big(I-A^{\frac{1}{2}}\overline{M}A^{\frac{1}{2}}\big)A^{\frac{1}{2}},$
    we obtain
[…]
    A straightforward analysis yields
    (4.13) $\|\mathbf{u}-\mathbf{u}^{(0)}\|_{E_{1}}^{2}\leq\big(1-\lambda_{n-r+1}(\widetilde{M}A)\big)\|\mathbf{u}-\mathbf{u}^{(0)}\|_{A}^{2},$
    where we have used the fact
    $\lambda(\overline{M}A)=\lambda(\widetilde{M}A).$
  • It is easy to see that
    $E_{2}=A^{\frac{1}{2}}\big(I-A^{\frac{1}{2}}M^{T}A^{\frac{1}{2}}\big)\Pi\big(I-A^{\frac{1}{2}}MA^{\frac{1}{2}}\big)A^{\frac{1}{2}}.$
    In light of (4.1), we have
    (4.14) $\|\mathbf{u}-\mathbf{u}^{(0)}\|_{E_{2}}^{2}\leq(1-\delta_{\rm TG})\|\mathbf{u}-\mathbf{u}^{(0)}\|_{A}^{2}.$
  • Note that
    $E_{1}-E_{2}$ $=(I-MA)^{T}A^{\frac{1}{2}}(I-\Pi)A^{\frac{1}{2}}(I-MA)$
    $=(I-MA)^{T}A^{\frac{1}{2}}(I-\Pi)^{T}(I-\Pi)A^{\frac{1}{2}}(I-MA)$
    $=(I-MA)^{T}(I-\Pi_{A})^{T}A(I-\Pi_{A})(I-MA)$
    $=E_{\rm TG}^{T}AE_{\rm TG}.$
    By (2.7) and (4.8), we have
    (4.15) $\|\mathbf{u}-\mathbf{u}^{(0)}\|_{E_{1}-E_{2}}^{2}=\|\mathbf{u}-\mathbf{u}_{\rm TG}\|_{A}^{2}\leq\rho^{2}\|\mathbf{u}-\mathbf{u}^{(0)}\|_{A}^{2}.$
According to (4.12), we deduce that
$\|\mathbf{u}-\mathbf{u}_{\rm ITG}\|_{A}^{2}\leq\|\mathbf{u}-\mathbf{u}^{(0)}\|_{E_{1}-E_{2}}^{2}+\varepsilon^{2}\|\mathbf{u}-\mathbf{u}^{(0)}\|_{E_{2}}^{2},$
which, combined with (4.14) and (4.15), leads to
(4.16) $\|\mathbf{u}-\mathbf{u}_{\rm ITG}\|_{A}\leq\sqrt{\rho^{2}+\varepsilon^{2}(1-\delta_{\rm TG})}\|\mathbf{u}-\mathbf{u}^{(0)}\|_{A}.$
In addition, we get from (4.12) that
$\|\mathbf{u}-\mathbf{u}_{\rm ITG}\|_{A}^{2}\leq(1-\varepsilon^{2})\|\mathbf{u}-\mathbf{u}^{(0)}\|_{E_{1}-E_{2}}^{2}+\varepsilon^{2}\|\mathbf{u}-\mathbf{u}^{(0)}\|_{E_{1}}^{2},$
which, together with (4.13) and (4.15), yields
(4.17) $\|\mathbf{u}-\mathbf{u}_{\rm ITG}\|_{A}\leq\sqrt{\rho^{2}+\varepsilon^{2}\big(1-\rho^{2}-\lambda_{n-r+1}(\widetilde{M}A)\big)}\|\mathbf{u}-\mathbf{u}^{(0)}\|_{A}.$
The desired estimate then follows immediately by combining (4.16) and (4.17). ∎
@@
Recall that
$\|\mathbf{u}-\mathbf{u}_{\rm TG}\|_{A}\leq\|E_{\rm TG}\|_{A}\|\mathbf{u}-\mathbf{u}^{(0)}\|_{A},$
where $\|E_{\rm TG}\|_{A}<1$. For the sake of practicality, we have made a more general assumption in Theorem 4.3. Another main assumption made in Theorem 4.3 is (4.9), which, in fact, characterizes the relative accuracy of the coarse solver $\mathscr{B}_{\rm c}\llbracket\cdot\rrbracket$.
@@
$\|\mathbf{u}_{\rm TG}-\mathbf{u}_{\rm ITG}\|_{A}=\|P(\mathbf{e}_{\rm c}-\hat{\mathbf{e}}_{\rm c})\|_{A}=\|\mathbf{e}_{\rm c}-\hat{\mathbf{e}}_{\rm c}\|_{A_{\rm c}},$